OpenStudy (anonymous):

I find my mistake, please help. The two curves y1 = bx^2 and y2 = ln x "touch" at a point. Calculate the exact value of b. I get that y1 = y2 (1) & y1' = y2' (2) I solve for b in (1) and substitute it in (2) But I end up getting 1/(2 sqrt(e)) The answer is: 1/(2e)

OpenStudy (anonymous):

I can't find*

OpenStudy (anonymous):

strange, i don't get either of those things i must be doing something wrong first one gives \[bx^2=\ln(x)\] \[b=\frac{\ln(x)}{x^2}\] second one gives \[2bx=\frac{1}{x}\] \[b=\frac{1}{2x^2}\]

OpenStudy (anonymous):

books answer is right though what did you compute?

OpenStudy (anonymous):

if we substitute (1) in (2)\[b(\frac{ 1 }{ 2 })=\frac{ 1 }{ 2 } \ln \frac{ 1 }{ 2b }\] \[b=\ln 1-\ln 2b\] \[b=-\ln 2b\] cant move from here

OpenStudy (anonymous):

i get \[\frac{1}{2x^2}=\frac{\ln(x)}{x^2}\] \[\ln(x)=\frac{1}{2}\] \[x=\sqrt{e}\] then find \(b\) by setting \[b(\sqrt{e})^2=\ln(\sqrt{e})\] \[eb=\frac{1}{2}\] \[b=\frac{1}{2e}\]

OpenStudy (anonymous):

i made mistake its \[\frac{ 1 }{ 2 }=\frac{ 1 }{ 2 } \ln \frac{ 1 }{ 2b }\] \[1=\ln \frac{ 1 }{ 2b }\] \[e^1=1/2b\] \[b=\frac{ 1 }{ 2e }\]